40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8908     Entry time: Tue Jul 23 16:39:31 2013
Author: Koji 
Type: Update 
Category: General 
Subject: Full IFO alignment recovered 

[Annnalisa Koji]

Full alignment of the IFO was recovered. The arms were locked with the green beams first, and then locked with the IR.

In order to use the ASS with lower power, C1:LSC-OUTPUT_MTRX_9_6 and C1:LSC-OUTPUT_MTRX_10_7 were reduced to 0.05.
This compensates the gain imbalance between TRX/Y siganls and the A2L component in the arm feedback signals.

Despite the IFO was aligned, we don't touch the OPLEVs and green beams to the vented IFO.

Attachment 1: alignment.png  18 kB  | Hide | Hide all
alignment.png
ELOG V3.1.3-